# On De Giorgi’s lemma for variational interpolants

### v2

@@ 4.2 Regularity of the variational interpolant
Let ${\mathcal{R}}:X\to[0,\infty)$ satisfy Hypothesis 3.1 and be radially differentiable. Then, ${\mathcal{R}}^{*}{\circ}\partial{\mathcal{R}}:X\to X^{*}$ fulfills (4.3) if and only if $\mathfrak{P}:X\to[0,\infty)$ is locally Lipschitz:
$\forall\,M>0\ \ \exists\,C_{M}>0\ \ \forall\,u_{1},u_{2}\in X\text{ with }\max_{i=1,2}\|u_{i}\|\leq M\colon\ |\mathfrak{P}(u_{1}){-}\mathfrak{P}(u_{2})|\leq C_{M}\|u_{1}{-}u_{2}\|\,.$ (4.8)
Proof. It suffices to observe that $\mathfrak{P}(u)={\mathcal{R}}(u)+({\mathcal{R}}^{*}{\circ}\partial{\mathcal{R}})(u_{1})$ and to and use that ${\mathcal{R}}$ is itself locally Lipschitz since ${\mathcal{R}}^{*}$ has superlinear growth at infinity.
As as immediate corollary of the above characterization, we have that property (4.3) is stable under the sum of dissipation potentials.
@@
Let ${\mathcal{R}}_{i}:X\to[0,\infty)$, $i=1,2$, satisfy Hypothesis 3.1 and be radially differentiable. If (4.3) is valid for ${\mathcal{R}}_{i}$, $i=1,2$, then it is valid for ${\mathcal{R}}_{1}+{\mathcal{R}}_{2}$ as well.
@@
Ultimately, $p$-homogeneous dissipation potentials (including the degenerate case $p=1$, which however falls outside the scope of this paper), provide examples for the validity of (4.3).
@@
Let ${\mathcal{R}}:X\to[0,\infty)$ be positively homogeneous of degree $p$ for some $p>1$. Then, ${\mathcal{R}}^{*}{\circ}\partial{\mathcal{R}}$ fulfills (4.3).
Proof. Firstly, we claim that
$\mathfrak{P}(v)=p{\mathcal{R}}(v)\qquad\text{for all }v\in X\,.$ (4.9)
To show (4.9), it suffices to observe that, by definition of $\partial{\mathcal{R}}(v)$ and $p$-homogeneity of ${\mathcal{R}}$, we have for all $\xi\in\partial{\mathcal{R}}(v)$
@@
Then, we divide the above estimate by $(\lambda{-}1)$ for all $\lambda>1$, and take the limit as $\lambda\to 1^{+}$, thus obtaining $p{\mathcal{R}}(v)\geq\mathfrak{P}(v)$. The converse inequality follows by dividing by $(\lambda{-}1)$ for all $10<\lambda<$ 1 and sending $\lambda\to 1^{-}$.
From (4.9) and the local Lipschitz continuity of ${\mathcal{R}}$ we deduce that $\mathfrak{P}$ satisfies (4.8), hence (4.3) holds thanks to Lemma 4.5.
@@
Remark 4.8 (Condition (4.3) in the framework of Example 3.2)
Getting back to Example 3.2, we see that a sufficient condition for the dissipation potential ${\mathcal{R}}(v)=\int_{\Omega}\mathrm{R}(v(x))\;\!\mathrm{d}x$ to satisfy property (4.3) is that the function $\mathrm{P}:{\mathbb{R}}\to[0,\infty)$, defined by $\mathrm{P}(r)=r\,\mathrm{R}^{\prime}(r)$ for $r\neq 0$ and by $\mathrm{P}(0)=0$ such that $\mathfrak{P}(v)=\int_{\Omega}\mathrm{P}(v(x))\;\!\mathrm{d}x$ (cf. (3.23)), is, itself, locally Lipschitz continuous.
4.3 De Giorgi’s estimate for general dissipation potentials
We now show that, if we drop the radial differentiability condition (3.17) on ${\mathcal{R}}$, De Giorgi’s estimate can be still retrieved for $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}})$ featuring general dissipation potentials ${\mathcal{R}}$. Our strategy consists in bringing into play the Yosida approximations $({\mathcal{R}}_{\eta})_{\eta}$ of ${\mathcal{R}}$, which clearly satisfy (3.17) guaranteeing the validity of De Giorgi’s identity for the corresponding $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}}_{\eta})$. We will then study the limit $\eta\to 0^{+}$ in (4.1b).
Remark 4.9
An approach based on Yosida regularization was also developed in [Bac21, Chap. 3] in order to extend the existence result for $\mathrm{gBGS}$ provided in [MRS13a], without the radial differentiability condition that was adopted therein (recall the discussion prior to Prop. 3.8). In [Bac21] the limit $\eta\to 0^{+}$ was performed on the level of the evolution equation. Here we resort to Yosida approximation to get rid of radial differentiability for the very proof of De Giorgi’s estimate. Hence, the limiting process $\eta\to 0^{+}$ will be more delicate than in [Bac21], because variational interpolants are not absolutely continuous in $[0,\sigma_{0}]$, in contrast to solutions of the regularized $\mathrm{gBGS}$.
For taking the limit as $\eta\to 0^{+}$, we will rely on the following closedness condition, tailored to the multi-valued operator $\partial_{\mathcal{R}}{\mathcal{E}}:X{\times}(0,\infty){\times}X\rightrightarrows X^{*}$ from (3.11).
Hypothesis 4.10 (Closedness of $\partial_{\mathcal{R}}{\mathcal{E}}$)
For every fixed $u^{o}\in\mathrm{dom}({\mathcal{E}})$ the conditioned subdifferential $\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\cdot;\cdot):(0,\infty){\times}X\rightrightarrows X^{*}$ is closed on energy sublevels, i.e.
$\forall\,E>0:\ \left\{\begin{array}[]{c}(\sigma_{n},u_{n},\xi_{n})\rightharpoonup(\sigma,u,\xi)\ \text{ in }(0,\infty){\times}X{\times}X^{*},\\ u_{n}\in S_{E}\text{ and }\xi_{n}\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma_{n};u_{n})\ \text{ for all }n\in{\mathbb{N}}\end{array}\right\}\ \Longrightarrow\ \xi\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;u)\,.$ (4.12)
Again, we highlight that, although condition (4.12) is required on sequences $(u_{n})_{n}$ weakly converging in $X$, some extra compactness may be derived from the information that the sequence $(u_{n})_{n}$ lies in the fixed energy sublevel $S_{E}$ and that the sequence $(\xi_{n})_{n}$ with $\xi_{n}\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;u_{n})$ is bounded.
We point out two crucial consequences of Hypothesis 4.10.
Lemma 4.11
Let the $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}})$ satisfy Hypotheses 3.1 and 4.10. Then,
  1. 1.
    $\mathcal{C}_{\mathcal{R}}(u^{o},\sigma;u):=\min\big{\{}\,{\mathcal{R}}^{*}({-}\xi)\>\big{|}\>\xi\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;u)\,\big{\}}\quad\text{for all }u\in J_{\sigma}(u^{o});$ (4.13)
  2. 2.
    For every measurable selection $(0,\infty)\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$ there exists a measurable selection
@@ 4.3 De Giorgi’s estimate for general dissipation potentials
    $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\mathfrak{A}_{{\mathcal{R}}}(\sigma,\widetilde{u}_{\sigma}):=\mathrm{argmin}\big{\{}\,{\mathcal{R}}^{*}({-}\xi)\>\big{|}\>\xi\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\widetilde{u}_{\sigma})\,\big{\}}$ (4.14a)
    such that
@@
    $\mathcal{C}_{\mathcal{R}}(u^{o},\sigma;u)={\mathcal{R}}^{*}({-}\widetilde{\xi}_{\sigma})\quad\text{for all }\sigma>0\,.$ (4.14b)
Proof. Property (4.13) is easily checked via the direct method by relying on (4.12). For (4.14b), observe that, as a consequence of Hypothesis 4.10, the multivalued mapping
$\mathfrak{A}_{{\mathcal{R}}}:(0,\infty){\times}X\rightrightarrows X^{*};\quad\mathfrak{A}_{{\mathcal{R}}}(\sigma,u):=\mathrm{argmin}\big{\{}\,{\mathcal{R}}^{*}({-}\xi)\>\big{|}\>\xi\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;u)\,\big{\}}$ (4.15)
is upper semicontinuous with respect to convergence in ${\mathbb{R}}$ for $\sigma$ and weak convergence for $u$. Therefore, for every measurable selection $(0,\infty)\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$ the multivalued mapping $\sigma\mapsto\mathfrak{A}_{{\mathcal{R}}}(\sigma,\widetilde{u}_{\sigma})$ is measurable, as it is given by the composition of measurable mappings. Then, [CaV77, Thm. 3.22] grants the existence of a measurable selection $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\mathfrak{A}_{{\mathcal{R}}}(\sigma,\widetilde{u}_{\sigma})$.
We are now in a position to state our second main result for Banach GS. The following result only states the existence of at least one measurable variational interpolant satisfying De Giorgi’s estimate. So far, we are not able to show that all measurable interpolants have this property.
Theorem 4.12 (De Giorgi’s estimate for general potentials)
Let $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}})$ satisfy Hypotheses 3.1 and 4.10, and let $\sigma_{*}>0$.
@@
Then, for every $u^{o}\in\mathrm{dom}({\mathcal{E}})$ there exists a measurable variational interpolant $(0,\infty)\ni\sigma\mapsto\widetilde{u}_{\sigma}\in X$ fulfilling De Giorgi’s estimate, and in particular there exists a measurable selection $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\widetilde{u}_{\sigma})$ along which there holds
@@
$\forall\,\sigma\in(0,\sigma_{*}]\colon\quad{\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\big{)}+\int_{0}^{\sigma}{\mathcal{R}}^{*}(-\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{o}).$ (4.16)
@@
Prior to carrying out the proof, let us introduce the main tool we are going to resort to: Since $X$ is reflexive, Asplund’s renorming theorem [Asp68], see also [BaP12, Thm. 1.105], ensures that there exists an equivalent norm on $X$ (still denoted by $\|\cdot\|_{X}$), such that, correspondingly, both $X$ and $X^{*}$ are strictly convex (namely, the mappings $x\mapsto\|x\|_{X}^{2}$ and $\xi\mapsto\|\xi\|_{X^{*}}^{2}$ are strictly convex). With this, the Moreau-Yosida approximations $({\mathcal{R}}_{\eta})_{\eta>0}$ of ${\mathcal{R}}$ are defined via
${\mathcal{R}}_{\eta}\colon X\to[0,\infty)\qquad{\mathcal{R}}_{\eta}(v):=\inf_{w\in X}\left(\frac{\|w{-}v\|^{2}}{2\eta}{+}{\mathcal{R}}(w)\right)\,.$
@@
    $\forall\,S>0\ \ \exists\,C_{S}>0\ \ \forall\,\eta\in(0,1]\ \ \forall\,v\in X:\quad{\mathcal{R}}_{\eta}(v)\leq S\ \ \Longrightarrow\\ \|v\|\leq C_{S}\,;$ (4.17)
@@
    the family $({\mathcal{R}}_{\eta})_{\eta}$ Mosco-converge to ${\mathcal{R}}$ as $\eta\downarrow 0$.

### v3

@@ 4.2 Regularity of the variational interpolant
Let ${\mathcal{R}}:X\to[0,\infty)$ satisfy Hypothesis 3.1, be radially differentiable, and suppose that $\mathcal{R}$$^{*}$ has superlinear growth at infinity, i,e.,
$\lim_{\|\xi\|_{*}\to\infty}\frac{{\mathcal{R}}^{*}(\xi)}{\|\xi\|_{*}}=\infty\,.$ (4.14)
Then, ${\mathcal{R}}^{*}{\circ}\partial{\mathcal{R}}:X\to X^{*}$ fulfills (4.9).
Proof. It follows from (4.14) that the subdifferential $\partial{\mathcal{R}}:X\rightrightarrows X^{*}$ is a bounded operator, cf. [Bre73, Prop. 2.14]. Then, (4.9) immediately follows, taking into account that $\mathcal{R}$$^{*}$ is itself Lipschitz on bounded sets since $\mathcal{R}$} has superlinear growth at infinity.
We also remark that, if $\mathcal{R}$$^{*}$ is superlinear, then property (4.9) is equivalent to the Lipschitz continuity on bounded sets of the mapping $\mathfrak{P}$ from (3.21). This follows from the fact that $\mathfrak{P}(u)={\mathcal{R}}(u)+({\mathcal{R}}^{*}{\circ}\partial{\mathcal{R}})(u)$, and $\mathcal{R}$} is Lipschitz continuous on bounded sets. Exploiting this observation, we can readily prove that property (4.9) is stable under the sum of dissipation potentials.
@@
Let ${\mathcal{R}}_{i}:X\to[0,\infty)$, $i=1,2$, fulfill Hypothesis 3.1 be radially differentiable, and satisfy (4.14). Then, ${\mathcal{R}}_{1}+{\mathcal{R}}_{2}$ satisfies (4.9).
@@
Ultimately, $p$-homogeneous dissipation potentials provide examples for the validity of (4.9).
@@
Let ${\mathcal{R}}:X\to[0,\infty)$ be positively homogeneous of degree $p$ for some $p>1$. Then,
$\mathfrak{P}(v)=p{\mathcal{R}}(v)\qquad\text{for all }v\in X$ (4.15)
and $\mathcal{R}$$^{*}{\circ}\partial{\mathcal{R}}$ fulfills (4.9).
Proof. First of all, we show (4.15) (which, in fact, also holds for $p=1$). Indeed, it suffices to observe that, by definition of $\partial{\mathcal{R}}(v)$ and $p$-homogeneity of $\mathcal{R}$}, we have for all $\xi\in\partial{\mathcal{R}}(v)$
@@
Then, we divide the above estimate by $(\lambda{-}1)$ for all $\lambda>1$, and take the limit as $\lambda\to 1^{+}$, thus obtaining $p{\mathcal{R}}(v)\geq\mathfrak{P}(v)$. The converse inequality follows by dividing by $(\lambda{-}1)$ for all $0<\lambda<1$ and sending $\lambda\to 1^{-}$.
From (4.15) and the fact that $\mathcal{R}$} is Lipschitz continuous on bounded sets, we deduce that $\mathfrak{P}$ satisfies (4.9).
@@
Example 4.10 (Example 3.5 re-revisited)
Getting back to Example 3.5, let us additionally assume that $\mathrm{R}$ is differentiable on ${\mathbb{R}}{\setminus}\{0\}$: then the dissipation potential ${\mathcal{R}}(v)=\int_{\Omega}\mathrm{R}(v(x))\;\!\mathrm{d}x$ is radially differentiable by Remark 3.14. Since we have required $\mathrm{R}^{*}$ to have $p^{\prime}$-growth with $p^{\prime}>1$, $\mathcal{R}$$^{*}$ has superlinear growth at infinity, and then Lemma 4.7 applies, ensuring that property (4.9) holds.
4.3 De Giorgi estimates for general dissipation potentials
We now return to the general case in which $\mathcal{R}$} is allowed to take the value $\infty$ with an open domain $\mathop{\mathrm{dom}}({\mathcal{R}})$ in $X$ (cf. Hypothesis 3.1). We will show that, if we drop the radial differentiability condition (3.20) on $\mathcal{R}$}, the De Giorgi simple and improved estimates can be still retrieved for $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}})$ featuring general dissipation potentials $\mathcal{R}$}, see Theorem 4.12 ahead.
Our strategy for proving both results consists in replacing $\mathcal{R}$} by the Yosida approximations $({\mathcal{R}}_{\eta})_{\eta}$, where we use a suitable norm on the reflexive Banach space $X$ such that all ${\mathcal{R}}_{\eta}$ are differentiable. Thus, for all $\eta>0$ we find measurable variational interpolants
@@ 4.3 De Giorgi estimates for general dissipation potentials
$\sigma\mapsto\widetilde{u}^{\eta}_{\sigma}\in\bm{J}_{\eta}(\sigma):=\mathrm{Argmin}\Big\{\>\sigma{\mathcal{R}}_{\eta}\big(\tfrac{\displaystyle 1}{\displaystyle\sigma}(u{-}u^{\circ})\big)+{\mathcal{E}}(u)\;\Big|\;u\in X\>\Big\}$ (4.16)
satisfying the De Giorgi identity for the corresponding $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}}_{\eta})$. We will then study the limit $\eta\to 0^{+}$.
A related approach based on Yosida regularization was developed in [Bac21, Chap. 3] in order to extend the existence result for $\mathrm{gBGS}$ provided in [MRS13], without the radial differentiability condition that was adopted therein (recall the discussion prior to Prop. 3.11). In [Bac21] the limit $\eta\to 0^{+}$ was performed on the level of the solutions of the evolution equation.
Here, we will resort to Yosida approximation to get rid of radial differentiability for the very proof of the De Giorgi estimate. Hence, the limiting process $\eta\to 0^{+}$ will be more delicate than in [Bac21], because variational interpolants are not absolutely continuous in $0,\sigma_{0}]$, in contrast to solutions of the regularized $\mathrm{gBGS}$. Nevertheless, starting from our approximations $(\sigma\mapsto\widetilde{u}^{\eta}_{\sigma})_{\eta>0}$ we will be able to find a limiting function $\sigma\mapsto\widetilde{u}_{\sigma}\in X$ that such that
@@
$\forall\,\sigma\in(0,\sigma_{*})\ \exists\,(\eta^{\sigma}_{k})_{k\in{\mathbb{N}}}:\quad\eta^{\sigma}_{k}\to 0\ \text{ and }\ \widetilde{u}^{\eta^{\sigma}_{k}}_{\sigma}\rightharpoonup\widetilde{u}_{\sigma}\quad\text{for }k\to\infty,$
see Proposition 4.15. This pointwise weak convergence along a $\sigma$-dependent subsequence will the crucial point to pass to the liminf, in the De Giorgi identity for $(X,{\mathcal{E}},{\mathcal{R}}_{\eta})$, to derive the De Giorgi estimates, also based on the Mosco-convergence as $\eta\to 0^{+}$ of the functionals
$\Phi_{\sigma}^{\eta}(u):=\sigma{\mathcal{R}}_{\eta}\big(\tfrac{\displaystyle 1}{\displaystyle\sigma}(u{-}u^{\circ})\big)+{\mathcal{E}}(u)$ (4.17)
involved in (4.16), to the functional $\Phi_{\sigma}$ for fixed $\sigma>0$, which means
$\left\{\begin{array}[]{ll}\forall\,(u_{\eta})_{\eta},u\in X\,:\ u_{\eta}\rightharpoonup u\ \Rightarrow\ \liminf_{\eta\to 0^{+}}\Phi_{\sigma}^{\eta}(u_{\eta})\geq\Phi_{\sigma}(u),\\ \forall\,u\in X\ \exists\,(\hat{u}_{\eta})_{\eta}\,:\hat{u}_{\eta}\to u\text{ and }\limsup_{\eta\to 0^{+}}\Phi_{\sigma}^{\eta}(\hat{u}_{\eta})\leq\Phi_{\sigma}(u).\end{array}\right$. (4.18)
This will follow easily from the Mosco-convergence of the potentials $({\mathcal{R}}_{\eta})_{\eta}$ stated in the upcoming Lemma 4.13.
For the proof of the simple De Giorgi estimate it will be sufficient to invoke the lower semicontinuity of the $\mathcal{R}$}-slope of $\mathcal{E}$}, namely $\mathcal{S}_{{\mathcal{R}}}$, in the limit passage in the De Giorgi identity holding for the regularized potentials. To obtain the improved the De Giorgi estimate, we will have to impose an additional condition on the lower semicontinuity of the conditioned $\mathcal{R}$}-slope, namely $\mathcal{C}_{{\mathcal{R}}}$:
Hypothesis 4.11 (Qualified lower semicontinuity of $\mathcal{C}_{{\mathcal{R}}}$)
Fix $\sigma\in(0,\sigma_{*})$ and $u^{\circ}\in\mathop{\mathrm{dom}}({\mathcal{E}})$ and consider $u_{\eta}\in\bm{J}_{\eta}(\sigma)$ for $\eta\in{]0,1[}$. Then,
$u_{\eta}\rightharpoonup u\quad\Longrightarrow\quad\liminf_{\eta\to 0}\mathcal{C}_{{\mathcal{R}}_{\eta}}(\sigma;u_{\eta})\geq\mathcal{C}_{{\mathcal{R}}}(\sigma;u).$ (4.19)
We will discuss sufficient conditions for this hypothesis after stating the main result of this subsection. We emphasize that the following result is weaker than the De Giorgi identity in two ways. First we only have the inequalities instead of the equality, and second we only prove the existence of at least one measurable interpolant, whereas the De Giorgi identity holds for all measurable interpolants with $\widetilde{u}_{\sigma}\in\bm{J}(\sigma)$. So far, we are not able to show that all measurable interpolants have this property.
Let us also mention in advance that, mirroring the statement of Theorem 4.2, under the additional Hypothesis 4.1, we are in a position to prove the existence of a measurable force selection for which the improved De Giorgi estimate holds. An analogous statement could be given for the simple De Giorgi estimate; we choose to omit it to avoid overburdening the exposition.
Theorem 4.12 (De Giorgi estimates for general $\mathcal{R}$})
@@
Let the $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}})$ satisfy Hypothesis 3.1, and let $\sigma_{*}>0$. Then, for every $u^{\circ}\in\mathrm{dom}({\mathcal{E}})$ there exists a measurable variational interpolant $\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}\subset X$ fulfilling the simple De Giorgi estimate
@@
$\forall\,\sigma\in(0,\sigma_{*})\colon\quad{\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)+\int_{0}^{\sigma}{\mathcal{S}}_{\mathcal{R}}(\widetilde{u}_{\sigma})\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{\circ}).$ (4.20)
f Hypothesis 4.11 holds, then this interpolant also satisfies the improved De Giorgi estimate
$\forall\,\sigma\in(0,\sigma_{*})\colon\quad{\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)+\int_{0}^{\sigma}\mathcal{C}_{{\mathcal{R}}}(\sigma;\widetilde{u}_{\sigma})\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{\circ}).$ (4.21a)
Thus, under the additional Hypothesis 4.1 there exists a measurable force selection $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma})$ satisfying the improved De Giorgi estimate with force selection
$\forall\,\sigma>0:\quad{\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)+\int_{0}^{\sigma}{\mathcal{R}}^{*}(-\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{\circ})\,.$ (4.21b)
We postpone the proof of this main result to Section 4.4.
@@
Prior to discussing the rather abstract lower semicontinuity property (4.19) on the conditioned slope $\mathcal{C}_{{\mathcal{R}}}$, we need to pin down some of the properties of the Moreau-Yosida approximations $({\mathcal{R}}_{\eta})_{\eta>0}$. First of all, let us settle their definition: Since $X$ is reflexive, Asplund’s renorming theorem [Asp68], see also [BaP12, Thm. 1.105], ensures that there exists an equivalent norm on $X$ (still denoted by $\|\cdot\|_{X}$), such that, correspondingly, both $X$ and $X^{*}$ are strictly convex (namely, the mappings $x\mapsto\|x\|_{X}^{2}$ and $\xi\mapsto\|\xi\|_{X^{*}}^{2}$ are strictly convex). With this, the Moreau-Yosida approximations $({\mathcal{R}}_{\eta})_{\eta>0}$ of $\mathcal{R}$} are defined via
${\mathcal{R}}_{\eta}\colon X\to[0,\infty)\qquad{\mathcal{R}}_{\eta}(v):=\inf_{w\in X}\left(\frac{\|w{-}v\|_{X}^{2}}{2\eta}{+}{\mathcal{R}}(w)\right)\,.$
@@
    $\forall\,S>0\ \ \exists\,C_{S}>0\ \ \forall\,\eta\in(0,1]\ \ \forall\,v\in X:\quad{\mathcal{R}}_{\eta}(v)\leq S\ \ \Longrightarrow\ \ \|v\|\leq C_{S}\,;$ (4.22)
@@
    the family $({\mathcal{R}}_{\eta})_{\eta}$ Mosco-converge to $\mathcal{R}$} as $\eta\downarrow 0$; in particular, we have
    $v_{\eta}\rightharpoonup v\ \Longrightarrow\ {\mathcal{R}}(v)\leq\liminf_{\eta\to 0}{\mathcal{R}}_{\eta}(v_{\eta})\quad\text{and}\quad\xi_{\eta}\rightharpoonup\xi\ \Longrightarrow\ {\mathcal{R}}^{*}(\xi)\leq\liminf_{\eta\to 0}{\mathcal{R}}^{*}_{\eta}(\xi_{\eta}).$
  3. 3.
    For all $(v_{\eta})_{\eta}\subset X$, $(\xi_{\eta})_{\eta}\subset X^{*}$
    $\left.\begin{array}[]{cl}v_{n}\to v&\text{as }\eta\to 0^{+},\\ \xi_{\eta}\rightharpoonup\xi&\text{as }\eta\to 0^{+},\\ \xi_{\eta}\in\partial{\mathcal{R}}_{\eta}(v_{\eta})&\text{for all }\eta>0\end{array}\right\}\quad\Longrightarrow\quad\xi\in\partial{\mathcal{R}}(v)\,.$ (4.23)